If 2sinQ-√3=0 find the the value of  2sinQ into cosQ

Grade:10

1 Answers

Rajat
4 years ago
2sinQ-√3=0
or, SinQ = √3/2
Thus Q = π/3
So cosQ= cosπ/3= ½
So, 2SinQ×CosQ=  2(√3/2)(1/2)= √3/2
Please upvote!!!